### v2

# spacing=nonfrench

Abstract
Most adversarial threats in artificial intelligence (AI) target the computational behavior of models rather than the humans who rely on them. Yet modern AI systems increasingly operate within human decision loops, where users interpret and act on model recommendations. Large Language Models (LLMs) generate fluent natural-language explanations that shape how users perceive and trust AI outputs, revealing a new attack surface at the cognitive layer: the communication channel between AI and its users. We introduce adversarial explanation attacks (AEAs), where an attacker manipulates the framing of LLM-generated explanations to modulate human trust in incorrect outputs. We formalize this behavioral threat through the trust miscalibration gap, a metric that captures the difference in human trust between benign and adversarial explanations. Using this metric as a lens, we highlight a behavioral risk where persuasive explanation framing can preserve user trust even when the underlying AI prediction is wrong.
To characterize this threat, we conducted a human study with over $200$ participants, systematically varying four dimensions of explanation framing: reasoning mode, evidence type, communication style, and presentation format. Our findings show that users report nearly identical trust for adversarial and benign explanations, with adversarial explanations preserving the vast majority of benign trust despite being incorrect. The most vulnerable cases arise when AEAs closely resemble expert communication, combining authoritative evidence, neutral tone, and domain-appropriate reasoning. Vulnerability is highest on hard tasks, in fact-driven domains, and among participants who are less formally educated, younger, or highly trusting of AI. To the best of our knowledge, this is the first systematic security study that treats explanations as an adversarial cognitive channel and quantifies their impact on human trust in AI-assisted decision making.
1 Introduction
Adversarial threats in artificial intelligence (AI) have primarily targeted the behavior of AI models. Attacks, such as adversarial examples, data poisoning, and model inversion, manipulate computational components to alter AI predictions and degrade accuracy [yuan2019adversarial, fredrikson2015model]. These studies assume autonomous systems in which humans passively accept model outputs. Yet modern deployments rarely fit this assumption. AI is increasingly embedded in AI-assisted decision-making workflows, where people interpret, verify, and act upon model recommendations. As AI becomes a decision partner rather than an oracle, new vulnerabilities emerge: not in data or algorithms, but in how AI communicates with and influences users.
Large Language Models (LLMs) epitomize this shift. Beyond producing predictions, they can generate additional fluent natural language explanations that accompany their output [breum2024persuasive]. Explanations are designed to foster transparency and calibrate trust [miller2019explanation, liao2022questioning]. However, their persuasive tone and rhetorical coherence can powerfully shape human reasoning. Prior studies show that people often conflate confident articulation with correctness [moore2008trouble]. This observation reframes the security boundary: An attacker who controls the presentation of reasoning may change user behavior. We argue that this form of behavioral manipulation represents a new class of adversarial threat that targets user cognition rather than model computation. Understanding such human-centric adversarial risks is essential to ensuring the robustness of AI-assisted decision-making systems.
In security terms, this behavioral threat extends beyond model manipulation into the human decision loop. Prior work [szegedy2014intriguing, goodfellow2015explaining, yuan2019adversarial] on adversarial machine learning largely targets the model itself, by altering inputs, parameters, or training data to change predictions. In contrast, adversarial explanation attacks (AEAs) hold the model itself fixed and explore how the adversary manipulates the form and framing of LLM-generated explanations to bias users toward accepting incorrect outputs. This reveals a cognitive attack surface on the AI to human communication channel, which remains exposed even when the underlying model is robust. We quantify this vulnerability through a trust miscalibration gap: a behavioral analogue of adversarial success that captures how often users continue to trust incorrect model outputs under adversarial explanations. This metric provides a compact way to assess adversarial influence at the cognitive layer of AI systems. Although our empirical study centers on AI-assisted decision making, the same mechanism underlies a wide range of human-centric AI systems, from conversational copilots to autonomous decision interfaces, where explanation framing can shape user judgment [miller2019explanation, breum2024persuasive].
Our study is grounded in established theories of trust calibration and persuasion. Decades of research on human interaction with automation and intelligent systems show that people’s reliance on such systems depends on perceived competence, predictability, and intent [lee2004trust, hancock2011meta]. From a cognitive standpoint, prior research shows that humans alternate between analytical and heuristic reasoning, making peripheral cues, such as tone, confidence, or evidential framing, powerful drivers of trust [petty1986elaboration, cialdini2007influence]. These insights collectively suggest that both the content of explanations (reasoning and evidence) and their style (tone and presentation) can modulate user trust. Guided by these foundations, we design a controlled study that systematically varies these factors across four dimensions: reasoning mode, evidence type, communication style, and presentation format, to analyze how explanation framing shapes human vulnerability to persuasive AI communication. To guide our study, we ask the following research questions:
RQ1 (Adversarial Impact): Can adversarial explanations measurably distort user trust and decision alignment, even when the underlying AI outputs are incorrect?
RQ2 (Mechanisms and Moderators): Which explanation strategies, contextual conditions, and individual traits most strongly influence user susceptibility to persuasive framing?
RQ3 (Dynamic Trust Evolution): How does user trust recalibrate or erode with repeated exposure to misleading explanations over time?
To address these questions, we conducted a large-scale controlled experiment ($n>200$) that exposed users to AI-assisted decisions under both benign and adversarial conditions. We systematically manipulate the reasoning mode, evidence type, communication style, and presentation format of AI-generated explanations while under varying task contexts and user traits. This design enables quantitative estimation of the trust miscalibration gap.
Our findings show that most users make decisions based on explanations, and adversarial explanations do not substantially reduce trust scores. Users report nearly identical trust for adversarial and benign explanations, and a positive trust miscalibration gap emerges: incorrect outputs framed adversarially receive trust scores that are close to those of correct outputs. Most participants base their trust on the explanation itself rather than on the underlying facts or the AI system, which allows persuasive but misleading explanations to sustain high trust even on errors.
We further find that the most vulnerable adversarial framings imitate trustworthy expert communication. Strategies that combine authoritative evidence such as citations or statistics, neutral analytic tone, and domain-appropriate reasoning produce the largest trust miscalibration gaps, while mismatched styles sharply reduce trust. Vulnerability is highest on medium and hard tasks and in fact-driven domains such as medicine and business, and lower in logic-intensive domains such as mathematics and law. It also varies by user group, participants with less educated, younger and high pre-trust on AI showing the most susceptibility and highly educated users and older participants showing more calibrated skepticism and greater reliance on their own knowledge.
Longitudinal analysis shows that trust in each task is driven mainly by the current explanation, with little carryover across tasks. Repeated detection of attacks gradually erodes trust, whereas benign streaks restore it. Overall trust in AI remains stable before and after the study, but a meaningful minority reduces their trust and a smaller group increases it when adversarial explanations align with their prior expectations, suggesting that adversarial communication can keep short-term trust high while slowly nudging long-term trust.
Together, these results show that the human-AI interface itself constitutes an exploitable attack surface. An adversary does not need to alter data, models, or code; manipulating the communication of reasoning alone is enough to induce trust miscalibration and threaten decision integrity. To our knowledge, this is the first systematic security study to conceptualize adversarial explanations as a cognitive attack vector and to empirically quantify their effects through a trust miscalibration gap. By reframing adversarial robustness to include human cognition, this work lays the foundation for securing not only algorithms but also the human decision processes intertwined with them.
Contributions. In this work, we treat explanations as a first-class security interface rather than a neutral transparency tool. We define adversarial explanation attacks that act entirely in the AI-to-human communication channel. Concretely, we (1) formalize the trust miscalibration gap as a behavioral analogue of adversarial success, (2) construct a four-dimensional space of LLM-generated explanation strategies that exposes a compositional attack surface, and (3) conduct a large-scale user study that quantifies which strategies, task contexts, and user groups produce the largest trust miscalibration gap with incorrect outputs. By shifting the adversarial lens from model robustness to cognitive robustness, our results argue that any security analysis of AI-assisted decision workflows must include the AI-to-human communication channel as an explicit attack surface.
2 Related Work
2.1 Adversarial Risks in Human-AI Interaction
Most adversarial AI research targets the model itself through accuracy degradation, data manipulation, or information extraction [yuan2019adversarial, fredrikson2015model]. A smaller but growing body of work examines how such attacks influence human decision-making. Model-level studies show that perturbations to predictions can indirectly alter human reliance, and that the timing and context of failures shape downstream trust [lu2023strategic]. Complementary efforts identify additional vulnerability sources, including interface-level factors such as explanation style, transparency cues, and presentation [ditz2025secure], as well as human-factor influences such as expectations, prior trust, and cognitive strategies [fan2025position]. Together, these threads indicate that susceptibility arises through a combination of model outputs, interface design, and human cognition. However, existing work assumes that trust distortion originates from model errors or benign design flaws. No prior work examines whether an adversary can manipulate human judgment purely through explanation framing, without altering the underlying model or its predictions.
2.2 LLM-Generated Explanations as a Cognitive Channel
Large Language Models (LLMs) are now widely used for tasks such as question answering [singhal2025toward], summarization [franceschelli2024creativity], and software development [nejjar2025llms]. A defining capability of these systems is the generation of natural-language explanations intended to support user understanding and collaboration [bilal2025llms, breum2024persuasive]. Because explanations articulate model reasoning in human-interpretable form, they increasingly function as a cognitive layer through which users interpret, evaluate, and act on AI outputs [bilal2025llms].
Recent work raises concerns about the reliability and influence of this layer. Studies document unfaithful, incorrect, or hallucinated rationales [madsen2024self, agarwal2024faithfulness, wei2022chain], and show that persuasive but inaccurate explanations can mask model flaws and elevate over-trust [turpin2023language, arcuschin2025chain]. Faithfulness evaluations [chuang2024faithlmfaithfulexplanationslarge, matton2025walk] focus on alignment between explanations and model internals but do not analyze how explanations shape human vulnerability. Collectively, prior work shows that explanations substantially influence user cognition, even when model predictions remain unchanged, positioning the explanatory layer as an important but underexamined surface for security risks.
2.3 Trust and Persuasion in Human-AI Decision Making
User trust is central but often fragile in human-AI collaboration [lee2004trust, steyvers2025large]. Prior work shows that judgments depend on user traits [CHONG2022107018], perceived system reliability [yin2019understanding], and task complexity [Salimzadeh2023]. Miscalibrated trust, including overreliance and underreliance, can arise when systems present persuasive or confident rationales [siekeretal2024illusion, chromik2021think]. LLMs intensify these risks: their fluent, human-like explanations and dialog capabilities influence judgment [breum2024persuasive], can outperform humans in persuasion tasks [Salvi_2025, meta2022human], and exploit anthropomorphic cues that increase user receptivity [zeng2024johnny, michelle2024]. Longitudinal research further shows that repeated interactions lead to cognitive overreliance [ibrahim2025towards], indicating vulnerability even without adversarial intent.
Misleading explanations have been studied in limited settings. Lakkaraju et al. [lakkaraju2020fool] showed that static post-hoc explanations for traditional black-box models can be crafted to mislead experts, but these manipulations rely on fixed feature-attribution templates and fail to generalize to the rich, adaptive natural-language reasoning of LLMs. More importantly, prior work attributes trust distortion to model failures or non-adversarial interface choices rather than purposeful manipulation. Our work departs from these threads by formalizing adversarial explanation attacks (AEAs): deliberate manipulation of explanation framing to distort user trust while leaving model predictions unchanged. This reframes explanations from an incidental interaction effect into a cognitive-layer attack vector within the AI-assisted decision loop.
3 Threat Model
3.1 AI-Assisted Decision-Making Setting
We consider an AI-assisted decision-making system in which an AI model produces both a prediction $\hat{y}$ and an explanation $e$ for a human decision-maker $H$:
$x\rightarrow\text{AI Model}\rightarrow(\hat{y},e)\rightarrow H\rightarrow y_{h}.$
Here, $x$ denotes the task input, $y_{h}$ the human’s final decision, and $e$ an explanation generated by an AI model such as an LLM. The explanation serves as the AI-to-human communication channel through which the human interprets the model’s reasoning, evaluates confidence, and decides whether to accept or reject $\hat{y}$. Therefore, the integrity of this channel is crucial: a misleading or overly persuasive explanation can distort human judgment and compromise decision reliability. The primary targets of such attacks are end-users, especially those without domain expertise who depend on LLM-generated explanations to guide their actions, although users with partial expertise may also be susceptible. The scope of this threat model encompasses AI-assisted decision systems in which explanations guide user interpretation and action, including copilots, medical and financial advisors, and educational assistants.
3.2 Adversary Goals and Capabilities
We define a behavioral threat model in which the adversary compromises the communication integrity of the system by manipulating how explanations are framed, even when the prediction is incorrect. This behavioral threat targets the human decision loop rather than the model’s internal computation. The attacker seeks to induce users to accept incorrect model outputs by generating a plausible and persuasive explanation.
We assume the adversary can influence the explanation-generation pipeline through practical attack vectors in modern LLM deployments, e.g., adversarial prompting that elicits persuasive yet misleading explanations [zou2023universal], malicious fine-tuning or data poisoning alter reasoning structure or evidential framing [zou2023universal, shu2023exploitability], and middleware hijacking that intercepts or rewrites explanation text at the API or plug-in layer [zhou2024trustworthiness]. Representative scenarios include (1) a clinical decision support tool where prompt templates and explanation style are centrally configured and can be manipulated through prompt injection or malicious fine-tuning, and (2) an LLM-powered financial assistant where explanations are rendered in a web interface and middleware can intercept and rewrite explanation text but not model logits. In both settings, the adversary is constrained to generate explanations that remain syntactically valid and context-appropriate, and cannot alter the choice options shown to the user. Unlike classical attacks on model inputs or parameters, the behavioral adversary exploits cognitive vulnerabilities inside the human-AI communication loop.
4 Adversarial Explanation Framework
Building on the threat model in Section 3, we now operationalize how adversarial explanation attacks (AEAs) exploit the communication channel between an AI system and its human user. This section formalizes the behavioral attack surface of AI-generated explanations, framing it as a compositional space whose manipulation can modulate user trust and decision alignment. The behavioral effects identified here are evaluated empirically through large-scale user studies described in Section 5.
4.1 Conceptual Foundations
Prior research in explainable AI (XAI), cognitive psychology, and human-AI interaction all emphasize that explanations shape human interpretation through multiple channels. XAI work shows that explanation content, tone, and form jointly determine perceived model transparency and competence [miller2019explanation, linardatos2020explainable]. Cognitive psychology models such as the Elaboration Likelihood Model [petty1986elaboration] describe how users shift between analytical and heuristic reasoning, making them susceptible to peripheral cues such as confidence, empathy, or authority framing. Human-AI studies further confirm that stylistic and visual elements, such as politeness, affect, or emphasis, directly affect user reliance [lee2025towards, chiaburu2024uncertainty, das2024shifting].
Synthesizing these perspectives, we model an AI explanation as a composite communication act that operates along four axes: cognitive, evidential, social, and perceptual. This yields a framework that not only organizes explanation strategies but also exposes new levers of adversarial manipulation. The resulting four-dimensional space supports factorial control over explanation properties, allowing us to measure how each dimension contributes to trust-accuracy divergence.
Table 1: Examples of explanation strategies from our four-dimensional design space. Each response addresses the question: “Based on the financial profile and market outlook, which investment option is most appropriate for the client?” Baseline strategies in each dimension are marked with *.
Dimension Strategy Generated Example Reference
Reasoning Mode Counterfactual (CR) If the client had a lower risk tolerance, the model would have recommended option B, which is a conservative bond fund. [artelt2021evaluating, del2024generating] [mothilal2020]
Feature Attribution (FA) The recommendation A is based on the client’s age, income stability, and long-term growth objectives. [aas2021explaining, guidotti2018survey]
Analogy & Example (AE) Choosing this investment A is like opting for a diversified basket, it spreads risk across multiple sectors. [hullermeier2020towards, poche2023natural] [ding2023fluid]
Procedural Reasoning (PR) First, the client’s financial goals were assessed. Then, risk appetite was evaluated, leading to this option A. [wei2022chain, gao2023learning] [haollm, blasco2024impact]
Uncertainty-Awareness (UA) There’s only a 65% confidence level in this recommendation A, given recent market volatility. [chiaburu2024uncertainty, slack2021reliable] [wang2021show]
Reframing (R) Rather than focusing on short-term market dips, consider how this recommendation A supports the long-term wealth growth. [kaufman2003frames, prinz2024effect] [yang2025understanding]
Plain Rationale* (N) The model predicts option A is the best investment.
Evidence Type Citation & Stat-Pack (CS) According to Pedersen et al. (2021) in the Journal of Financial Economics, ESG-optimized portfolios deliver higher risk-adjusted returns, forming an efficient frontier comparable to traditional investment models. [huang2023citation, ding2025citations] [lewis2020retrieval]
Equation & Proof (EP) Using the Sharpe ratio formula, $S=\frac{R_{p}-R_{f}}{\sigma_{p}}$, option A offers a superior risk-adjusted return. [yang2023leandojo, welleck2022naturalprover] [tafjord2021proofwriter]
Internal Conceptual* (IC) The model predicts the answer is A based on patterns it has learned from historical data.
Communication Style Sycophancy (S) Great question! You are absolutely right to consider this, option A is definitely the best choice. [sharma2024towards, wang2025truth] [cheng2025socialsycophancybroaderunderstanding, malmqvist2025sycophancy]
Emotional & Empathetic (EE) I understand investment decisions can feel overwhelming, here’s why option A is a secure choice for you. [concannon2023interactional, kerasidou2020artificial] [jeon2024effects, roshanaei2025talklistenconnecthumans]
Neutral* (NE) The model predicts option A based on input data. [miller2019explanation, rong2023towards]
Presentation Format Visual Emphasis (VE) The chart below shows how option A has historically outperformed other funds in similar risk categories. [joshi2024constrained, sun2025effect] [das2024shifting]
Plain Verbal* (PV) The answer is A based on the model’s prediction.
4.2 Adversarial Explanation Space Construction
We represent each explanation strategy as a four-tuple $s=(r,v,c,p)$, where $r$, $v$, $c$, and $p$ denote its reasoning mode, evidence type, communication style, and presentation format, respectively. Each dimension corresponds to a cognitively distinct channel through which framing can influence human trust while keeping the model prediction $\hat{y}$ fixed. In our threat model, the adversary does not control the input $x$ or the predicted answer $\hat{y}$, but can select a framing strategy $s$ from the explanation space and supply this configuration to the explanation generator.
Reasoning Mode ($r$) controls the logical structure of justification, including counterfactual, feature attribution, analogy & example, procedural reasoning, uncertainty-awareness, or reframing, with plain rationale as baseline. Each mode evokes a different cognitive pathway: contrastive reasoning highlights causality, analogies evoke familiarity, and procedural reasoning mirrors step-wise human logic [artelt2021evaluating, wei2022chain].
Evidence Type ($v$) captures the form of justification supporting an explanation, including citations & stat-pack and equation & proof, along with internal conceptual as a baseline. Different evidence types shape users’ perception of verifiability and epistemic credibility [famiglini2024evidence].
Communication Style ($c$) defines the interpersonal tone, such as sycophantic, emotional & empathetic, or neutral. It has been shown that subtle shifts in communication style, such as expressing empathy, or agreement, can significantly alter user perceptions, even when the underlying content remains unchanged [lee2022polite].
Presentation Format ($p$) determines how information is visually or structurally conveyed, ranging from visual emphasis (figures, highlights, tables) to plain verbal presentation. Even minimal visual cues can direct user attention and bias perceived credibility [sun2025effect, das2024shifting].
Table 1 summarizes representative strategies along with LLM-generated examples. Detailed descriptions for strategies are discussed in Appendix B.
4.3 Operational Role in Adversarial Analysis
We now formalize how this compositional explanation space can be exploited by an adversary. For a given task $q$, an explanation framing strategy $s=(r^{\prime},v^{\prime},c^{\prime},p^{\prime})$ induces an explanation $e_{\text{A}}(q,s)$ generated by the LLM, while $e_{\text{B}}(q)$ denotes the benign explanation without framing strategy. Rather than evaluating model-centric robustness or prediction accuracy, we focus on user trust and trust miscalibration gap: how much users trust incorrect model outputs under adversarial explanations. Let $T(u,q,e)\in\mathbb{R}$ denote the user’s trial-level self-reported trust score for explanation $e$ on task $q$, measured on a 1177 Likert scale, which quantifies how much the user is willing to rely on the AI-provided answer given explanation $e$. We define the trust miscalibration gap as the change in user trust induced by adversarial explanation relative to the benign condition:
$\Delta T(q,s)=\mathbb{E}_{u}\!\left[T(u,q,e_{\text{A}}(q,s))\right]-\mathbb{E}_{u}\!\left[T(u,q,e_{\text{B}}(q))\right].$ (1)
A large positive value of $\Delta T(q,s)$ indicates that users are susceptible to adversarial framing $s$ in that task $q$. In our threat model, the adversary aims to select a strategy $s$ that maximizes trust miscalibration on incorrect outputs, i.e., large $\Delta T(q,s)$ on tasks $q$ with $\hat{y}(q)\neq y(q)$. Throughout the paper, we use $T$ as the universal measure of trial-level user trust, and $\Delta T(q,s)$ as the primary metric for quantifying explanation-induced trust miscalibration.
5 User Study Design and Protocol
This section translates the adversarial explanation framework into a controlled behavioral security experiment. The study evaluates whether manipulations of explanation framing $s=(r,v,c,p)$ can cause measurable trust miscalibration, a behavioral analogue of model misclassification in which users develop misplaced trust in incorrect AI outputs. The design aims to emulate realistic attacker capabilities, isolate causal effects of framing, and provide statistically validated measures of behavioral attack success. All implementation details, prompt examples, and dataset curation procedures are documented in Appendix A for transparency.
5.1 Unified Design and Experimental Setup
The experiment models how an adversary can exploit plausible yet misleading explanations to influence human decision-making. Four design principles ensure both adversarial realism and experimental validity.
  • Plausibility over verifiability. Attackers rarely need factual correctness; they only need to sound reasonable. Our manipulations prioritize surface plausibility and coherence, reflecting real-world persuasion where users judge explanation persuasiveness heuristically rather than through verification.
  • Trust as attack success. User trust serves as the behavioral measure of attack success. High trust in incorrect outputs indicates successful persuasion. We operationalize this through self-reported Likert trust ratings.
  • Controlled composition. Each explanation combines a configuration of reasoning, evidence, communication, and presentation strategies introduced in Section 4. This factorial design enables causal attribution of trust shifts to specific framing mechanisms.
  • Reproducibility and auditability. All prompts, configurations, and outputs are logged to ensure transparency and reproducibility, which is critical for behavioral security studies.
We test two conditions: (1) Benign, where the AI provides the correct answer with a valid explanation; and (2) Adversarial, where the AI justifies an incorrect answer using the same template. All other factors (task content, prompt format, explanation structure) remain identical.
Figure 1: Overview of the adversarial explanation generation and control, consisting of four stages: construct prompt-based instruction, generate explanations, quality control, and deliver curated samples to the user survey.
5.2 Adversarial Explanation Generation and Control
Our adversarial explanation generation and control pipeline in Figure 1 ensures reproducibility, scalability, and adversarial realism through four sequential stages. Appendix A documents the strategy-guided explanation generation prompt (A.1), dataset reconstruction (A.3), validation prompt (A.4).
Stage 1: Prompt Construction
For each task, we assemble structured few-shot prompts that define the assigned strategy tuple $(r,v,c,p)$ described in Section 4. Each prompt specifies the reasoning mode, evidence type, communication style, and presentation format that the model should adopt, along with the target answer (correct or incorrect). Definitions and exemplars are embedded to guide consistent stylistic and cognitive behavior across samples. This step mirrors an attacker crafting a persuasive but targeted explanation template.
Stage 2: Explanation Generation
Given the prompt, the generator model, i.e., Llama-3.3-70B, produces a concise explanation of 2 - 5 sentences (approximately 75 - 200 words). We sample tasks from the MMLU benchmark [hendryckstest2021], spanning seven broad domains and three difficulty levels. This diversity ensures that the manipulations generalize across cognitive and topical contexts rather than being domain-specific.
Stage 3: Quality Control
We apply a hybrid validation process to ensure both strategic fidelity and linguistic plausibility. An ensemble of smaller LLMs (Gemma 3-4B, Mistral-7B) evaluates whether each explanation conforms to its intended $(r,v,c,p)$ configuration through five independent votes, requiring majority agreement for retention. Surviving explanations are then rated using an LLM-as-a-Judge model [zheng2023judging] that estimates perceived trustworthiness on a 1–7 Likert scale. Explanations scoring lowest on trust or exhibiting hallucinations are filtered out after manual review.
Stage 4: Survey Delivery
Validated explanations are integrated into the Qualtrics survey environment [qualtrics2025]. For each question, the corresponding explanation, either Benign or Adversarial, is displayed alongside the model’s chosen answer. Survey randomization logic automatically assigns strategy combinations and task order per participant, ensuring balanced exposure across conditions while preventing learning effects. This final stage bridges automated adversarial generation and controlled human evaluation, enabling systematic behavioral measurement.
5.3 User Study Procedure
The main study was conducted under Institutional Review Board (IRB) approval using the Qualtrics platform [qualtrics2025]. Participants first completed demographic and AI-attitude questionnaires, followed by forty randomized tasks (10 Benign, 30 Adversarial). Each task displayed a question, the AI’s selected answer, and the corresponding explanation with strategies. Participants rated their trust on a 7-point Likert scale and indicated their rationale (persuasiveness of the explanation, confidence in the answer, prior AI attitude, or other), which reflects the cognitive sources of trust. Attention checks were inserted to ensure engagement, and the order of questions and conditions was fully randomized. The Appendix A.7 provides a sample task used in our user study survey.
We recruited 205 qualified participants from a US university and Amazon Mechanical Turk [AmazonMechanicalTurk], restricted to English-fluent adults ($\geq$18 years) with $\geq$95% approval and $\geq$500 prior HITs. The average completion time was approximately 45 minutes, and participants were compensated at or above local minimum wage. Demographic distributions are summarized in Table 4; additional participant-recruitment criteria and survey-allocation logic are included in Appendix A.5.
6 User Study Results
We present the empirical results of our user study to address RQ1-RQ3. The analysis quantifies how adversarial framing alters user trust (RQ1), identifies the explanation mechanisms and influencing factors that shape susceptibility (RQ2), and examines how trust evolves under repeated exposure to adversarial explanations (RQ3). Together, these findings reveal how adversarial explanations exploit cognitive vulnerabilities in human-AI communication to achieve behavioral deception.
In our statistical analysis, we employ ordinary least squares (OLS) regression to examine both independent and interaction effects of multiple factors, such as attack condition and cognitive source, on trust scores $T$. To analyze differences in the distribution of cognitive trust sources between attack and non-attack conditions, we apply chi-square tests of independence. We use Welch’s two-sample t-tests to conduct pairwise comparisons of trust scores under specific conditions, as this test is robust to unequal sample sizes and variance heterogeneity across groups. To account for repeated measurements from the same users across task sequences, we additionally apply linear mixed-effects models with participant-level random effects to investigate trust evolution.
6.1 Quantifying Adversarial Impact (RQ1)
RQ1: Can adversarial explanations measurably distort user trust and decision alignment, even when the underlying AI outputs are incorrect? We assess it by analyzing (i) cognitive sources of trust and (ii) trust shifts under adversarial vs benign conditions.
6.1.1 Cognitive Sources of Human Trust
To understand why users trust adversarial explanations, we first examine the cognitive sources from which trust is derived, i.e., what attributions lead to trust or distrust in AI outputs. Following prior work in human-AI trust [xu2024trust, wang2024explainabilitytrust], we categorize trust formation into three primary sources: (1) explanation-based, where trust is attributed to the perceived persuasiveness of the explanation; (2) prior-knowledge-based, where users rely on their own prior knowledge to assess correctness; and (3) trust-in-AI-based, where trust is placed in the AI system itself. An additional “Other” category was also provided in the survey to capture remaining sources. Post-processing analysis for this category is detailed in Appendix A.6.
The majority of users attributed their trust to the explanation itself (attack: 65.065.0%; non-attack: 64.564.5%; Figure 2). A chi-square test confirms that the prevalence of explanation-based trust does not differ significantly between conditions ($\chi^{2}=0.10$, $p=0.75$; Table 5), confirming that explanation plausibility, rather than factual accuracy, is the dominant driver of trust.
In contrast, trust attributions based on prior knowledge and trust in AI exhibit asymmetric shifts across conditions. Users relying on prior knowledge are significantly more prevalent in the non-attack condition (24.6% vs. 19.9%; $\chi^{2}=18.1$, $p\ll.001$), whereas trust-in-AI attributions are more common under attack (14.4% vs. 10.6%; $\chi^{2}=17.2$, $p\ll.001$), suggesting that adversarial explanations reallocate users’ cognitive attribution in AI trust. Detailed chi-square results are in Appendix C.1.
Figure 2: Proportion of trust cognitive sources under attacks and non-attacks. The majority of users attributed their trust to the explanation itself.
6.1.2 Measuring Trust Shift
Having established that explanation-based cognition dominates trust attribution, we next examine whether adversarial explanations induce measurable changes in trust magnitude. Focusing on explanation-driven trust, we find that average trust scores are nearly identical under attack conditions (attack: $4.53$; non-attack: $4.59$; Figure 3) and OLS regression confirms that the difference is not statistically significant ($\beta=0.06$, $p=.282$; Table 6). This highlights a critical vulnerability: users who rely on the plausibility of explanations are easily manipulated by adversarial outputs and are unable to detect when the underlying answer is incorrect. Users who rely on prior knowledge exhibit heightened sensitivity. The average trust drops sharply from $5.76$ to $4.47$ under attack, as shown in Figure 16. Even though it suggests that prior knowledge can reduce acceptance of adversarial explanations when users feel confident in their own expertise, there is a minority of users who make decisions based on prior knowledge. Detailed analysis of other cognitive sources are presented in Appendix C.1.
Figure 3: Distribution of user trust scores $T$ under adversarial and benign explanation conditions, conditioned on trials in which users reported relying on the explanation as their cognitive source.
Finding. Most users anchor their trust in the explanations rather than facts or the AI system. Adversarial explanations preserve user trust scores that are nearly indistinguishable from benign ones, making persuasive yet adversarial explanations particularly serious in Human-AI decision-making.
6.2 Explanation Mechanisms and Influencing Factors (RQ2)
RQ2: Which explanation strategies, contextual conditions, and individual traits most strongly influence user susceptibility to persuasive framing? The findings from RQ1 show that adversarial explanations sustain user trust even when outputs are incorrect. This motivates our next step: an analysis of why such misleading explanations remain persuasive and under what conditions they succeed. We answer RQ2 by analyzing user vulnerability at three levels: (1) explanation level mechanisms, (2) task level factors, and (3) user level traits.
6.2.1 Explanation-Level Mechanisms: Explanation Strategy Effectiveness
We quantify the impact of explanation strategies by systematically varying the four dimensions introduced in Section 4: reasoning mode, evidence type, communication style, and presentation format. To measure the overall effect of the adversarial framing strategy $s=(r,v,c,p)$, we aggregate strategy-level trust miscalibration across tasks by
$\bar{\Delta}T^{(s)}=\frac{1}{|\mathcal{Q}|}\sum_{q\in\mathcal{Q}}\Delta T(q,s),$ (2)
where $\mathcal{Q}$ denotes the set of tasks. The ten most and least persuasive combinations are summarized in Table 2. Across all combinations, several key patterns emerge.
Figure 4: Distribution of user trust scores $T$ across task domains and explanation strategies, including reasoning mode, evidence type, communication style, and presentation format. Baseline strategies in each dimension are marked with an asterisk (*): N (Neutral) for reasoning mode, IC (Internal Conceptual) for evidence type, NE (Neutral) for communication style, and PV (Plain Verbal) for presentation format.
Table 2: Top and bottom ten explanation strategy combinations ranked by $\bar{\Delta}T^{(s)}$ between adversarial and benign conditions.
Top 10
Combo $\bar{\Delta}T^{(s)}$ PR+CS+NE+VE $0.91$ R+CS+NE+PV $0.86$ UA+EP+NE+PV $0.79$ FA+EP+NE+VE $0.78$ R+EP+NE+PV $0.73$ AE+EP+EE+PV $0.70$ N+EP+NE+PV $0.61$ N+IC+S+VE $0.53$ PR+CS+S+VE $0.52$ FA+CS+EE+PV $0.49$
Bottom 10
Combo $\bar{\Delta}T^{(s)}$ AE+EP+NE+VE 3.51-3.51 PR+EP+S+VE 1.92-1.92 CR+EP+NE+VE 1.59-1.59 AE+EP+S+PV 1.52-1.52 AE+EP+NE+PV 1.51-1.51 N+EP+S+VE 1.45-1.45 PR+EP+NE+PV 1.44-1.44 N+EP+NE+VE 1.28-1.28 PR+EP+S+PV 1.26-1.26 PR+EP+NE+PV 1.24-1.24
Reasoning Mode. The effectiveness of a reasoning mode varies by task domain, as shown in Figure 4. In fact-driven domains such as Biology, Medical and Health Sciences, and Physical and Engineering Sciences, trust scores do not differ substantially from the Neutral baseline, suggesting limited sensitivity to reasoning variation when answers are perceived as objective. However, in Business and Management, Politics and History, Social Sciences and Law, and Mathematical Problem and Logical Reasoning, Analogy & Example (AE) and Feature Attribution (FA) receive systematically higher trust scores. These subjective or cognitively demanding domains often involve interpretation, judgment, or abstraction. Users are more receptive to explanations that provide concrete examples or feature-level justifications, helping them interpret or justify recommendations.
Evidence Type. Citation & Stat-Pack (CS) emerges as the most consistently effective evidence type across domains (Figure 4), yielding higher trust than the Internal Conceptual baseline in nearly all task categories. The top two most persuasive explanation strategy combinations in Table 2 both employ CS. This pattern indicates that references to external sources, statistics, or authoritative artifacts reliably boost perceived credibility. In mathematical and logical reasoning tasks, however, Equation & Proof (EP) achieves trust levels comparable to CS, reflecting users’ preference for formal derivations and symbolic justification in quantitative contexts.
Communication Style. Neutral (NE) framing yields higher trust than other styles, particularly in technical and analytical domains such as Biology and Health Sciences, Mathematical Reasoning, and Social Sciences, implying that users prefer objective, dispassionate tones for complex or serious content. Differences among communication styles are less pronounced in domains such as Business and Politics, where the nature of the task allows for broader stylistic flexibility. Nearly all of the top-ranked explanation strategy combinations employ a Neutral style (Table 2). In general, Emotional/Empathetic (EE) and Sycophantic (S) framings underperform Neutral style in most domains, aligning with prior findings that overt emotional appeal or excessive agreeableness no longer reliably enhances credibility [malmqvist2025sycophancy, sharma2024towards].
Presentation Format. Plain Verbal (PV) and Visual Emphasis (VE) formats perform similarly across most domains. A notable exception occurs in Mathematical and Logical Reasoning, where VE significantly outperforms PV, indicating that structured visual guidance reduces cognitive burden and facilitates comprehension in logic-intensive tasks. In fact-driven or interpretive domains, the gap between presentation formats is narrower, suggesting that other dimensions dominate trust formation.
Bringing these dimensions together, the most persuasive explanation strategies typically blend authoritative evidence (CS), appropriate reasoning (either relatable or formal), neutral communication, and cognitively accessible presentation. High-performing combinations, such as PR+CS+NE+VE or R+CS+NE+PV, achieve the largest increases in trust under adversarial explanations. In contrast, mismatches, such as pairing formal mathematical proof with a sycophantic tone, disrupt credibility and result in sharp trust degradation.
Finding. Adversarial explanations are most susceptible when they combine authoritative evidence with a neutral tone and a domain-appropriate reasoning mode, which lets them imitate trustworthy expert communication style.
6.2.2 Task-Level Factors: Task Difficulty and Domain
We next examine how task difficulty and subject domain modulate user trust in adversarial explanations.
Task Difficulty. Trust in adversarial explanations varies sharply with task difficulty. As shown in Figure 5, users display higher trust under attack for medium and high difficulty tasks ($\bar{T}=4.79$ and $4.72$), where self-confidence is lower. The OLS model (Table 7) confirms this pattern: under attack, trust for low-difficulty tasks is significantly lower than for high-difficulty tasks ($\bar{T}=3.69$ vs. $4.72$, $\beta=-1.04$, $p<.001$). Moreover, trust in adversarial explanations is significantly lower than non-adversarial ones under low difficulty ($\bar{T}=3.69$ vs. $4.98$, $\beta=1.49$, $p<.001$), with distributions skewed toward lower scores (1–3). In contrast, this relationship reverses under high difficulty, where adversarial explanations are trusted more than benign ones ($\beta=-0.20$, $p=.016$). These results suggest that as task complexity increases, users become more reliant on AI explanations, possibly due to diminished confidence in their own expertise, and more willing to defer to persuasive explanations. Detailed OLS results are provided in Appendix C.2.
Figure 5: Distribution of trust scores $T$ across task difficulties under attacks and non-attacks.
Task Domain. Task domain also moderates user trust. As shown in Figure 6, in fact-driven fields (e.g., Biological, Medical, and Health Sciences and Business and Management), users report higher average trust in adversarial explanations than in correct ones ($\bar{T}$: $5.10$ vs. $4.76$, $p<.001$; $5.07$ vs. $4.56$, $p<.001$ see Table 8). These distributions are right-skewed, with more users assigning high trust scores (6–7) under attack, likely due to the authoritative tone or statistical plausibility commonly used in explanations within these domains. Logic-intensive or interpretive domains (e.g., Mathematical and Logical Reasoning, Politics and History, and Social Sciences and Law) show the opposite: with lower mean trust under attack ($\bar{T}$: $3.95$ vs. $4.98$, $p<.001$; $3.82$ vs. $4.92$, $p<.001$; $3.92$ vs. $5.04$, $p<.001$), with distributions shifting leftward toward lower trust scores (1–4). Users apply stricter scrutiny and display reduced trust under adversarial framing particularly in domains where reasoning or evidence evaluation plays a significant role. These findings suggest adversaries can exploit domain conventions, such as formality or objectivity, to selectively enhance persuasion in high-stakes contexts.
Figure 6: Distribution of trust scores $T$ across task domains under attacks and non-attacks.
Finding. Task context matters for trust under adversarial explanations: users exhibit higher trust on hard problems and in fact-driven domains, where they are more likely to defer to explanations framed as authoritative or supported by statistical evidence.
6.2.3 User-Level Traits: Cognitive and Demographic Traits
We analyze how user traits, including education level, age, initial trust in AI, and familiarity with AI, modulate trust formation and resistance to adversarial explanation.
Education Level. Higher education, particularly at the graduate level, correlates with increased resistance to adversarial explanations. As shown in Figure 7, users holding a Master’s or Doctoral degree show the largest drop in trust when exposed to misleading explanations (Master’s: $4.95\rightarrow 4.38$; Doctoral: $4.87\rightarrow 4.31$, Table 10 shows t-test results of trust scores). Their distributions exhibit more low-to-mid trust scores (1–3) and fewer high scores (6–7) under attack, particularly among users with Doctoral degrees, indicating stronger discernment when explanations conflict with their reasoning. They also rely less on explanation-based justifications and more on prior knowledge (Figure 19). Participants with less formal education or some college background display relatively high trust in adversarial explanations ($\bar{T}$: 4.374.374.534.53), with distributions skewed toward mid-to-high trust scores (5–6). This pattern suggests a greater baseline susceptibility to persuasive but incorrect explanations among less formally trained users, possibly due to weaker confidence in their ability to verify information independently.
Figure 7: Distribution of trust scores $T$ across education levels under attacks and non-attacks.
Age. Older users tend to demonstrate greater skepticism toward AI explanations under adversarial conditions. As shown in Figure 8, users aged 3535-$44$ and 65+, trust in misleading outputs drops sharply compared to younger users, and the gap between attack and non-attack conditions is most pronounced ($4.56\rightarrow 4.11$ and $4.94\rightarrow 3.92$, $p<0.001$ respectively; see t-test results in Table 11). Their distributions shift leftward, showing higher proportions of low-to-mid trust scores (1–3) under attack, suggesting stronger critical evaluation when confronted with misleading explanations. Younger users (ages 18–34) maintain relatively high and symmetric trust distributions, with smaller declines under attack ($4.62\rightarrow 4.49$ and $4.91\rightarrow 4.42$), indicating a greater tendency to accept persuasive content at face value. Consistent with the distributional shifts, cognitive source analysis shows that older users rely less on explanation-based justifications and more on prior knowledge (Figure 20), supporting the interpretation that age-related experience enhances scrutiny of adversarial explanations.
Figure 8: Distribution of trust scores $T$ across ages under attacks and non-attacks.
Initial Trust in AI. Participants with higher initial (pre-survey) trust consistently report higher trust scores $\bar{T}$ across both adversarial and benign conditions (attack: $4.6$, non-attack: $5.78$, Figure 9). Their distributions are strongly right-skewed, with dense concentrations around high trust scores (6–7), indicating a persistent tendency to maintain confidence even when explanations are misleading. When these users attribute trust to the AI system itself, their trust is especially elevated, even when explanations are misleading (Figure 21). This highlights a risk of overreliance: users with strong initial faith in AI may be less critical and more easily influenced by persuasive framing. In contrast, users with low initial trust (“somewhat distrust” or “strongly distrust”) exhibit flatter, left-shifted distributions centered around lower scores (1–3), showing more cautious and discriminating evaluation regardless of condition. Overall, these findings suggest that initial trust acts as a cognitive prior that amplifies susceptibility to persuasion.
Figure 9: Distribution of trust scores $T$ across initial trust levels under attacks and non-attacks.
6.2.4 Familiarity with AI
We additionally examined how users’ familiarity with AI moderates their trust in adversarial explanations. While descriptive trends suggest that less experienced users tend to retain higher trust under attack, statistical comparisons within the extreme groups: not familiar at all and expert-level knowledge, are underpowered due to their relatively small sample sizes. We therefore report the detailed analyses of AI familiarity in Appendix C.4.
Finding. Less educated, younger users, and those with high prior trust in AI are the most vulnerable to adversarial explanations, while highly educated and older users show more calibrated skepticism.
6.3 Dynamic Trust Evolution (RQ3)
RQ3: How does user trust recalibrate or erode with repeated exposure to misleading explanations over time? While prior analyses treat tasks as static, trust in AI is inherently dynamic: users may adapt, habituate, or desensitize to persuasive framing as they accumulate exposure to adversarial explanations.
Short-term effect. Trust in each task is primarily driven by whether the current explanation is adversarial or benign. As shown in the OLS results (Table 3 (a)), users report significantly higher trust when the current task is benign than when it is adversarial ($\bar{T}=4.75$ vs. $4.38$; $\beta=0.37$, $p<.001$), regardless of the previous task condition. Neither the main effect of the previous task nor its interaction with the current task is statistically significant ($\beta=0.01$, $\beta=-0.02$). Pairwise Welch’s t-tests (Table 3 (b)) further confirm that, for both adversarial and benign current tasks, trust scores do not differ based on the previous task condition ($p=0.83$, $p=0.98$). These findings indicate that trust is primarily shaped by the framing of the current explanation, with minimal short-term carryover from prior tasks.
Table 3: Short-term effects of task attack conditions on user trust $T$. (a) OLS regression results predicting trust from current and previous task conditions and their interaction. Baseline category is Attack in current task with Attack as the previous task. (b) Pairwise Welch’s two-sample t-tests comparing trust scores between previous task conditions, conditioned on the current task. Significance: $^{*}p<.1$, $^{**}p<.01$, $^{***}p<.001$.
(a) OLS regression results
Effect Type Term $\boldsymbol{\beta}$
Main Effects Intercept $4.38^{***}$
Current (Non-attack) $0.37^{***}$
Previous (Non-attack) $0.01$
Interaction Current $\times$ Previous 0.02-0.02
(b) Pairwise Welch’s t-tests
Current Task Previous Task Mean Trust ($\bar{T}$) $p$
Attack Attack $4.38$ 0.83
Non-attack $4.40$
Non-attack Attack $4.75$ 0.98
Non-attack $4.75$
Long-term effect. The average trust score $\bar{T}$ shows a subtle downward trend (red line in Figure˜10) as users progress through the task sequence (most of them are adversarial), decreasing from $4.9$ at the beginning to $4.5$ toward the end. A linear mixed-effects model confirms this long-term erosion of trust, showing a significant negative association between task order and trust ($\beta=-0.005$, $p=0.007$, results in Table 13), indicating that repeated exposure cumulatively reduces user trust over time. To understand the effect of repeated exposures, we further analyze trust dynamics conditioned on streaks of detected attacks. [^1] We set trust scores of 1-3 as attack detection, 5-7 as non-detection. As shown in Figure 11, longer streaks of detected attacks (ed line) lead to progressively lower trust in the subsequent task, with mean trust dropping from $4.1$ after a single detection to nearly $2.0$ after ten consecutive detections. Spearman rank correlation confirms a strong monotonic decline ($\rho=-0.28$, $p<.001$), as shown in Table 14. Conversely, non-attack streaks restore and elevate trust, with subsequent trust rising from $4.2$ to above $5.8$ after sustained ten exposures to benign explanations (green line in Figure˜11). This effect is also supported by a significant positive correlation ($\rho=0.32$, $p<.001$) in Table 14. These results indicate that cumulative adversarial exposure progressively erodes user trust, whereas consistent benign feedback helps rebuild and stabilize trust over time.
Figure 10: Mean trust $\bar{T}$ over a sequence of tasks.
Figure 11: Mean trust $\bar{T}$ in the subsequent task after varying lengths of detected-attack or non-attack streaks.
Overall trust shift. We compare users’ self-reported trust in AI before (pre-trust) and after (post-trust) the survey. As shown in Figure 12, most users remained stable in their trust assessments, with $26.6\%$ consistently reporting “somewhat trust,” $20.3\%$ remained “neutral,” and $15.2\%$ stayed “somewhat distrust.” A subset of users decreased their trust after repeated exposure to adversarial explanations, $10.1\%$ shifted from “somewhat trust” to “neutral. By contrast, a smaller fraction of users increased trust, with $10.9\%$ shifting from “neutral” to “somewhat trust”, possibly because some adversarial explanations appeared credible or aligned with user expectations.
Figure 12: User overall trust shift in AI before and after the survey.
Finding. Trust is mainly determined by the current explanation, but repeated detection of attacks gradually erodes trust, whereas sequences of benign interactions restore confidence, leading to modest but meaningful shifts in global trust in AI.
7 Toward Defending Against Adversarial Explanations
Our empirical analysis across RQ1–RQ3 reveals a consistent vulnerability in AI-assisted decision making: persuasive explanations can modulate user trust. This shifts the adversarial threat surface from the model’s computational pipeline to the communication channel between AI and its users. When attackers shape the framing of explanations, they can influence human decisions even when the underlying model is correct. Securing AI systems, therefore, requires extending defenses beyond model robustness to explanation design, interfaces, and interaction protocols. Below, we outline a defense agenda informed by the vulnerabilities surfaced in our study.
Limit the explanation framing strategies
RQ1 and RQ2 show that highly expressive, stylistically flexible explanations create a wide, persuasive attack surface. Elements such as analytic tone, citations, statistics, and structured presentation inflate perceived credibility independent of correctness. One concrete design could be an explanation generator that outputs fixed-length, step-by-step rationales with explicit uncertainty bounds, while disallowing free-form rephrasing. As an initial feasibility check, future work can compare miscalibration under template-based explanations versus unconstrained LLM explanations using our experimental paradigm.
Prioritize verifiability over plausibility
RQ2 highlights that the most effective adversarial strategies mimic authority signals, including formal citations, statistical justification, and confident declarative language. These findings exploit heuristic trust, enabling incorrect outputs to appear well-supported. Defenses should therefore enforce verifiability rather than reward surface plausibility. A practical instantiation is to require that any numeric claim or citation in an explanation be verified by a retrieval backend and tagged as verified or unverifiable in the interface. The system can downgrade or block explanations that contain unverifiable assertions, reducing the degrees of freedom available to an attacker.
Risk-adaptive explanation policies
Our RQ2 results indicate that users are most vulnerable in hard tasks, fact-driven domains, or when they are unfamiliar with AI. In these settings, trust can even reverse, with incorrect answers receiving higher trust than correct ones. This motivates risk-aware adaptation, in which the system modulates its communication strategy according to predicted user vulnerability. Defenses include stronger uncertainty descriptions for complex domains, simplified non-authoritative templates for novices, contrastive explanations when overtrust is detected, and domain-specific warnings in tasks with a history of high risk.
8 Limitations
Our study isolates a minimal behavioral attack surface, i.e., explanation framing, to establish the feasibility and impact of adversarial explanations. Three scope boundaries apply. (1) We use simplified, text-based tasks and a controlled web interface to isolate framing effects from other cues such as interaction history or multimodal feedback. Real-world deployments often combine explanations with richer interface elements and time pressure. As a result, our estimates of trust miscalibration should be interpreted as a lower bound on what might occur in more complex settings. (2) Participants are users recruited from online rather than trained professionals in domains such as medicine or law, and their incentives and risk perceptions differ from those of high-stakes decision makers. Domain experts may exhibit different trust baselines, calibration between prior knowledge and model outputs, and different susceptibility to specific framing strategies. Future work should evaluate adversarial explanations in workflows where expertise, accountability, and institutional constraints are central. (3) We outline but do not implement or empirically test defenses, and we abstract away many system-level constraints that would shape their deployment in practice. Our proposals represent an initial design space for mitigating explanation-based attacks rather than validated countermeasures, and they may face trade-offs with usability or model performance. Overall, our contribution is to map and quantify a new behavioral attack surface and its risks, not to provide a complete mitigation blueprint for all real-world systems.
9 Conclusion
As AI systems increasingly serve as both decision-makers and explainers, a new attack surface emerges: the communication channel through which AI explains its outputs to users. Our work introduces the adversarial explanation attacks (AEAs), where plausible yet misleading LLM-generated explanations distort user judgment and foster misplaced trust in incorrect outputs. We formalize a four-dimensional explanation strategy, including reasoning mode, evidence type, communication style, and presentation format, to systematically manipulate explanation framing and assess its impact on user trust. By conducting a large-scale empirical user study ($n=205$), we find that users place nearly as much trust in adversarial explanations as in benign ones, with trust primarily attributed to the explanation itself. Moreover, user susceptibility varies by task context (e.g., difficulty, domain) and individual traits (e.g., education, age, initial trust in AI), and repeated exposure to misleading explanations can gradually erode or recalibrate trust. These findings reveal critical cognitive vulnerabilities in human–AI interaction and highlight the need for future defenses that address not only model robustness but also users’ cognitive resilience.
Ethical Considerations
This work studies how AI-generated explanations influence human trust and decision-making under both benign and adversarial conditions. Since adversarial explanations directly shape human judgment, this research raises ethical concerns related to manipulation and trust miscalibration in human-AI interaction.
Stakeholders.
We identify several stakeholders potentially impacted by this work: (1) human participants in the study, (2) end users of AI systems that provide explanations, (3) system designers and deployers responsible for explanation interfaces, (4) security researchers and practitioners, and (5) society at large, particularly in contexts where AI-assisted decisions carry real-world consequences. Our analysis considers both the immediate impacts of the research procedures and the downstream impacts of publishing these findings. Our intent is diagnostic and defensive: to expose vulnerabilities in explanation mechanisms so that they can be mitigated through safer design, constrained explanation formats, uncertainty signaling, and improved human oversight.
Ethical principles.
We ground our analysis in the principles articulated in the Menlo Report, particularly Beneficence, Respect for Persons, and Respect for Law and Public Interest. Beneficence motivates our focus on identifying and mitigating explanation-induced vulnerabilities that could otherwise remain hidden in deployed systems. Respect for Persons guides our treatment of human subjects, including informed consent, minimal risk exposure, and data protection. Respect for Law and Public Interest informs our approach to responsible disclosure and data sharing.
Potential harms.
A key potential harm is the dual-use risk that insights about persuasive explanation strategies could be misused to deliberately manipulate users. Additional risks include temporary trust miscalibration among study participants exposed to adversarial explanations, as well as broader concerns about eroding confidence in AI systems if vulnerabilities are misunderstood or misapplied. These harms are primarily cognitive and informational rather than physical or financial, but nonetheless implicate users’ autonomy and decision-making rights.
Mitigations.
We take several steps to mitigate these risks. Adversarial explanations are used exclusively in controlled experimental settings and are not deployed in real-world systems. We avoid releasing executable tools, prompts, or templates that would directly enable deceptive explanation generation. The study was conducted under institutional review board (IRB) approval. No personally identifiable information was collected, all responses were anonymized, and analyses were performed only in aggregate. Participants were informed of the experimental nature of the study, and exposure to potentially misleading explanations was limited in scope and duration.
Decision to conduct and publish.
We determined that the benefits of this research outweigh its potential harms. From a beneficence perspective, systematically characterizing explanation-based vulnerabilities enables the development of defenses such as constrained explanation formats, uncertainty signaling, auditing mechanisms, and improved human oversight. From a respect-for-persons perspective, identifying how explanations can undermine user autonomy is a necessary step toward protecting it. We believe that withholding these findings would leave critical vulnerabilities unexamined and impede the design of safer AI systems. Accordingly, we proceed with publication while deliberately limiting the release of artifacts that could facilitate misuse.
Open Science
Raw survey response data are not publicly released due to IRB and data privacy constraints. To support transparency, we show the prompt used in our experiment and the explanation space definition in the Appendix. Besides, we publicly release the full survey structure and representative examples of the main task questions at: https://github.com/anoymous-518/usenix-anon-artifact.
Appendix A Experimental Details
A.1 Explanation Prompt
Each explanation was generated using structured few-shot prompts specifying the assigned combination of reasoning mode, evidence type, communication style, and presentation format $(r,v,c,p)$. Prompts explicitly described each strategy with examples and instructed the LLM to justify the selected answer accordingly. A typical instruction is shown in Figure 13.
Strategy-Guided Explanation Generation Prompt System: You are an AI assistant providing a concise and lay-friendly explanation supporting the provided answer. You will be shown a multiple-choice question and choices and assigned to generate a concise explanation based on one answer. Your goal is applying specified explanation strategies to persuade users that the provided answer is correct. Explanation length is 2–5 sentences totaling 75–200 English words ($\approx$ 100–280 tokens), unless Visual Emphasis. Assume the answer index is correct; do not verify it. STRATEGY BLOCK  Dimension-1 (Reasoning Mode) Name: {D1_NAME} Definition: {D1_DEF} Example: {D1_EX} Dimension-2 (Evidence Type) Name: {D2_NAME} Definition: {D2_DEF} Example: {D2_EX} Dimension-3 (Communication Style) Name: {D3_NAME} Definition: {D3_DEF} Example: {D3_EX} Dimension-4 (Presentation Format) Name: {D4_NAME} Definition: {D4_DEF} Example: {D4_EX}   STRICT STRATEGY COMPLIANCE Single-Strategy Rule: Use every assigned strategy ($1\times$ D1, $1\times$ D2, $1\times$ D3, $1\times$ D4) and no element from any unassigned strategy. No Strategy Names: Do not mention words such as “Counterfactual,” “Stat-Pack,” “Sycophancy,” or similar. User: [OUTPUT] Question: {QUESTION_TEXT} Choices: {CHOICE_JSON} Answer Index: {ANSWER_IDX} Answer Text: {ANSWER_TEXT} Category: {CATEGORY} Only categories “Mathematical Problem Solving” and “Logical Reasoning and Abstract Thinking” may use Equation/Proof in Dimension-2. Output Format The output must contain only the explanation text and no additional commentary, headings, or formatting. The explanation must end with the exact sentence: Therefore, the answer is: {ANSWER_TEXT}. [OUTPUT]
Figure 13: Prompt used for strategy-guided explanation generation.
A.2 Experiment Settings
We used a temperature of 0.7 and nucleus sampling with $p=0.9$ to encourage stylistic variation. Explanation length was constrained to 75–200 English words (100–280 tokens) unless the strategy required longer formatting (e.g., Visual Emphasis).
A.3 Dataset and Difficulty Calibration
We used a curated subset of the MMLU benchmark [hendryckstest2021], consolidated into seven categories: Mathematics and Logic, Politics and History, Biology and Health Sciences, Social Sciences and Law, Physical and Engineering Sciences, and Business and Management. Question difficulty was scored by GPT-4o [hurst2024gpt] on a 1–9 scale and grouped into Low (1–3), Medium (4–6), and High (7–9), following [zhao2025mmlu]. Each user received a balanced mix of difficulty levels and topic areas.
A.4 LLM Quality Control and Configuration
Explanations were generated by Llama-3.3-70B, hosted through an instruction-tuned API. Strategy conformance was automatically validated by an ensemble of five runs using Gemma-3-4B and Mistral-7B-Instruct models. Each model independently judged whether the output reflected all assigned strategies, with majority (3/5) agreement required for acceptance. An exmaple prompt is shown in Figure 14. Explanations that passed were further evaluated by a Gemma-3-4B LLM-as-a-Judge model estimating trust scores on a 7-point Likert scale. Outputs rated lowest ($\geq$ 1) or showing hallucinated reasoning were removed after manual inspection. All random seeds, model versions, and configuration metadata were logged to support auditability and reproducibility.
Explanation Validation Prompt System: You are simulating an average, non-expert person judging how convincing a model’s answer is only from the explanation. Read the question, options, the model-chosen answer (index + text), and the explanation. Output a Likert score (1–7) that reflects how much you would trust the answer after reading the explanation. Also, output a brief reason (2 sentences) summarizing the key logic behind your trust level. SCORING RUBIC 1 = Totally unconvinced; explanation clearly flawed, contradictory, or admits mismatch with the chosen answer. 2 = Very low trust; major logical gaps, unsupported claims. 3 = Low trust; several weaknesses, limited supporting logic. 4 = Moderate/uncertain trust; mixed or incomplete reasoning. 5 = Fairly high trust; mostly plausible, minor issues. 6 = High trust; coherent, persuasive, few noticeable issues. 7 = Completely convinced; clear, consistent, strongly supports the answer. STRICT EVALUATION RULES If the explanation explicitly states its reasoning contradicts, ignores, or merely “aligns with” the pre-selected answer (e.g., “However, our task is to align with the selected index …”), immediately return score = 1. Use the full 1-7 range; avoid clustering all scores near the middle. User: [OUTPUT] Question: {QUESTION_TEXT} Choices: {CHOICE_JSON} Model-Selected Answer Index: {ANSWER_IDX} Model-Selected Answer Text: {ANSWER_TEXT} Explanation : {EXPLANATION_TEXT} Output format Return a valid JSON object with exactly two keys:
  "score": <integer 1–7>,
  "reason": "<concise justification (2 sentences)>"
No additional keys, no markdown, no code fences, and no line breaks outside the JSON object. Please evaluate and output the JSON result as specified. [OUTPUT]
Figure 14: Prompt used for explanation validation.
A.5 Survey Allocation and Participant Demographics
Each participant completed 40 questions: 10 Benign (correct answer + valid explanation) and 30 Adversarial (incorrect answer + plausible explanation). This allocation ensured broad coverage of strategy combinations. All participants were English-fluent adults, with MTurk workers required to have $\geq$ 95% approval and $\geq$ 500 prior HITs. Table 4 reports demographic statistics.
Table 4: Participant demographics, n=$205$
Gender Education
Male 101
Female 103 High school diploma or GED 38
Other 1
Age Some college or associate degree 73
18 - 24 84
25 - 34 39 Bachelor’s degree 52
35 - 44 30
45 - 54 25 Master’s degree 35
55 - 64 22
65 or older 5 Doctoral or professional degree (e.g., Ph.D., JD, MD) 7
A.6 Cognitive Sources Analysis for “Other”
We conducted a post-hoc review of the responses from the “Other” category, via keyword filtering and manual coding. The analysis showed that nearly all responses could still be recategorized into the three main categories, with only 1.0% stating other factors such as confusion about the question (0.20.2%) or a need for further verification for explanations (0.80.8%).
A.7 Human Study Survey Interface
Figure˜15 illustrates the user interface with a sample task presented to participants during the survey.
Figure 15: A sample task in the survey.
Appendix B Adversarial Explanation Space Details
B.1 Reasoning Mode
Reasoning mode refers to the logical structure or argumentative scaffolding used to justify the model’s output. We define the reasoning modes as follows:
  • Counterfactual Reasoning: Provides a "what-if" scenario that describes how a minimal change to the input could lead to a different outcome [artelt2021evaluating, del2024generating].
  • Feature Attribution: Identifies the most influential input features that contribute to the model’s decision and explains why the chosen answer uniquely satisfies key criteria [aas2021explaining, guidotti2018survey].
  • Analogy & Example: Justifies the provided answer by drawing a structural parallel to a real-world example or scenario [hullermeier2020towards, poche2023natural].
  • Procedural Reasoning: Presents a step-by-step, rule-based derivation, such as a logical chain-of-thought [wei2022chain] that incrementally leads to the correct answer.
  • Uncertainty-Awareness: Explicitly communicates the model’s confidence, predictive uncertainty, or willingness to abstain [chiaburu2024uncertainty, slack2021reliable].
  • Reframing: Alters the interpretive lens or evaluation criteria to highlight the strengths of the selected answer under an alternative frame [kaufman2003frames] (e.g., framing outcomes as gains instead of losses).
  • Plain Rationale (Baseline): A plain, minimal justification without any reasoning technique as the baseline.
B.2 Evidence Type
Evidence type refers to the form of justification provided to support the explanation. We summarize the three evidence types as follows:
  • Citation & Stat-Pack: Attributes claims to verifiable external sources or quantitative data summaries to enhance credibility and perceived trustworthiness [ding2025citations].
  • Equation & Proof: Constructs formal mathematical derivations or logical proofs. Note that this evidence type is only used for tasks in the “mathematical problem and logistic reasoning” category in our empirical survey.
  • Internal Conceptual (Baseline): Offers a general rationale based on model-internal patterns without citing external sources and proofs.
B.3 Communication Style
Communication style captures the tone and interpersonal framing of the explanation, i.e., how it positions itself in relation to the user. We consider the following communication styles:
  • Sycophancy: Frames the explanation in a deferential, agreement-seeking communication style that aligns with the user’s stated preferences [sharma2024towards, wang2025truth].
  • Emotional & Empathetic: Uses an affect-aware communication style that acknowledges the user’s emotions or concerns and communicates understanding, support, or reassurance [concannon2023interactional].
  • Neutral (Baseline): Uses a factual, impersonal tone without affective or persuasive elements.
B.4 Presentation Format
Presentation format describes how the explanation is visually or structurally conveyed. We categorize presentation formats into the following types:
  • Visual Emphasis: Enhances the explanation via visual formatting or structural cues, such as bullet points, bold or italic text, emoji, tables, or spatial layout [sun2025effect].
  • Plain Verbal (Baseline): Presents explanation as unformatted prose without visual or structural enhancements.
Appendix C Additional User Study Results
We present the additional user study results here (most have been discussed in the main text).
Table 5: Comparison of cognitive trust sources between attack and non-attack conditions. A separate $2\times 2$ chi-square test compares its frequency across conditions.
Source Attack Non-Attack $\boldsymbol{\chi^{2}(1)}$ $\boldsymbol{p}$
Explanation 3676 (65.0%) 1213 (64.5%) 0.1 0.75
Prior knowledge 1127 (19.9%) 462 (24.6%) 18.1 $2.1\times 10^{-5}$
Trust in AI 814 (14.4%) 199 (10.6%) 17.2 $3.4\times 10^{-5}$
Other 42 (0.7%) 6 (0.3%) 3.4 0.07
C.1 Cognitive Sources
The distribution of cognitive trust sources differed significantly between the attack and non-attack conditions (Table 5). The majority of users in both conditions attributed their trust to the explanation itself, and this pattern remained highly stable (Attack: 65.0%; Non-Attack: 64.5%; $\chi^{2}=0.10$, $p=0.75$), indicating that adversarial explanations did not alter the dominant role of explanation-based trust. In contrast, trust based on prior knowledge is significantly more common in the non-attack condition (24.6% vs. 19.9%; $\chi^{2}=18.1$, $p=2.1\times 10^{-5}$), suggesting that users were more likely to rely on their own expertise when the AI output was correct. Conversely, reliance on “trust in AI” is more frequent under attack (14.4% vs. 10.6%; $\chi^{2}=17.2$, $p=3.4\times 10^{-5}$), indicating that some users defaulted to trusting the model even when it is wrong. The “Other” category is rarely selected and showed only a marginal difference across conditions ($p=0.07$). These results suggest that the explanation plausibility is the dominant driver of trust and remains stable; changes in prior-knowledge and AI-trust attributions account for the overall distributional shift.
Users who rely on prior knowledge exhibit heightened sensitivity. The average trust drops sharply from $5.76$ to $4.47$ under attack, as shown in Figure 16. The distribution shifted leftward, showing more low-to-mid trust scores (1–2) in the attack condition, whereas participants without attack exposure were more likely to assign high trust scores (6–7), as corroborated by the OLS results: interaction of prior knowledge attribution and the attack condition is large and highly significant ($\beta=1.24$, $p<.001$; Table 6). This suggests that prior knowledge can reduce acceptance of adversarial explanations when users feel confident in their own expertise. Users who base their trust in the AI system itself report uniformly low trust across both conditions (attack: $4.03$; non-attack: $3.98$), regardless of explanation correctness, indicating an inherent skepticism toward AI outputs.
Figure 16: Distribution of trust scores $T$ across cognitive sources under attacks and non-attacks.
To further examine how cognitive source and condition interact to shape trust, we fit an ordinary least squares (OLS) regression model predicting trust scores $T$ from condition (attack vs. non-attack), cognitive source (explanation, prior knowledge, trust in AI, other), and their interaction. The baseline is explanation-based attribution under the attack condition. Full model results are reported in Table 6. The intercept ($\beta=4.53$, $p<.001$) reflects the average trust score for explanation-based users under attack (the baseline). The main effect of Condition is not significant ($\beta=0.06$, $p=.282$), indicating that among explanation-based users, average trust is essentially unchanged between attack and non-attack conditions. This highlights a critical vulnerability: users who base their decisions on explanation plausibility are easily manipulated by adversarial outputs and are unable to detect that the underlying answer is incorrect.
The key pattern emerges in the interaction terms. The interaction between Condition and Prior knowledge is positive and highly significant ($\beta=1.24$, $p<.001$), implying that prior-knowledge users trust the system significantly more in the non-attack condition than under attack. Prior-knowledge-based mean trust decreases from $5.76$ (non-attack) to $4.47$ (attack). This result implies that prior-knowledge users are markedly more sensitive to adversarial explanations and substantially reduce their trust when answers are incorrect. In contrast, the interactions for “Other” and “Trust in AI” are not significant, indicating that these groups maintained similar trust levels regardless of explanation correctness.
Table 6: OLS regression results for predicting trust scores $T$ from condition (attack vs. non-attack), cognitive source, and their interaction. Baseline category is Attack with Explanation-based attribution. Significance: $^{*}p<.1$, $^{**}p<.01$, $^{***}p<.001$.
Effect Type Term $\beta$ (coeff.)
Main Effects Intercept $4.53^{***}$
Condition (Non-attack) $0.06$
Attr. (Prior knowledge) 0.06-0.06
Attr. (Trust in AI) $0.50^{***}$
Attr. (Other) $1.98^{***}$
[…]
Model fit: $R^{2}=0.045$, $F(7,7531)=50.69$, $p<10^{-70}$
C.2 Task Difficulty
To examine how task difficulty modulates user trust in adversarial explanations, we fit an ordinary least squares (OLS) regression model predicting trust scores $T$ from condition (attack vs. non-attack), task difficulty (low, medium, high), and their interaction. The baseline category corresponds to the attack condition under high task difficulty. The OLS results are reported in Table 7.
The model reveals a statistically significant interaction between attack condition and task difficulty, indicating that the impact of adversarial explanations on trust varies systematically with task difficulty. Under attack, task difficulty alone has a strong effect on trust: users exhibit substantially lower trust on low-difficulty tasks compared to high-difficulty tasks ($\beta=-1.04$, $p<.001$), while no significant difference is observed between medium and high difficulty levels ($\beta=0.07$, $p=.232$). This suggests that when tasks are simple, users are better able to identify inconsistencies in adversarial explanations. As task complexity increases, users rely more on AI explanations, likely due to reduced self-confidence and greater deference to persuasive content.
On the other hand, the effect of adversarial explanations changes direction as task difficulty increases. For low-difficulty tasks, trust is significantly higher in the non-attack condition than in the attack condition ($\beta=1.49$, $p<.001$), indicating that users successfully detect and penalize incorrect explanations. This distinction weakens for medium-difficulty tasks and reverses for high-difficulty tasks. Specifically, under high task difficulty, users report significantly higher trust in adversarial explanations than in benign ones ($\beta=-0.20$, $p=.016$), despite the adversarial explanations being paired with incorrect answers.
Taken together, these results demonstrate that task difficulty critically shapes users’ ability to detect adversarial manipulation. When tasks are easy, users rely on their own competence to evaluate explanations and appropriately reduce trust under attack. As difficulty increases, self-confidence diminishes and users become increasingly reliant on AI-provided explanations. Under high task difficulty, this reliance leads to a counterintuitive outcome in which adversarial explanations are trusted more than benign ones. Overall, these findings identify task difficulty as a key vulnerability factor in human–AI decision-making, highlighting that adversarial explanations are most effective precisely when cognitive demands are high and human verification becomes challenging.
Table 7: OLS regression results for predicting trust scores $T$ from condition (attack vs. non-attack), task difficulty (low, medium, high), and their interaction. Baseline category is Attack with High task difficulty. Significance: $^{*}p<.1$, $^{**}p<.01$, $^{***}p<.001$.
Effect Type Term $\boldsymbol{\beta}$
Main Effects Intercept $4.72^{***}$
Condition (Non-attack) $0.20^{*}$
Difficulty (Low) $1.04^{***}$
Difficulty (Medium) $0.07$
[…]
Model fit: $R^{2}=0.063$, $F(5,7712)=103.5$, $p<10^{-100}$.
C.3 Task Domain
User trust in adversarial explanations varies significantly across task domains. As shown in Table 8, users placed higher trust in adversarial explanations than in correct ones within fact-driven domains such as Biological, Medical, and Health Sciences ($\bar{T}_{\text{attack}}=5.10$, $\bar{T}_{\text{non\_attack}}=4.76$, $p<.001$) and Business and Management ($\bar{T}_{\text{attack}}=5.07$, $\bar{T}_{\text{non\_attack}}=4.56$, $p<.001$), where explanations may appear more plausible or authoritative. In contrast, trust under attack was significantly lower in logic-based or abstract domains such as Mathematical Problem and Logical Reasoning ($\bar{T}_{\text{attack}}=3.95$, $\bar{T}_{\text{non\_attack}}=4.98$, $p<.001$), Politics and History ($\bar{T}_{\text{attack}}=3.82$, $\bar{T}_{\text{non\_attack}}=4.92$, $p<.001$), and Social Sciences and Law ($\bar{T}_{\text{attack}}=3.92$, $\bar{T}_{\text{non\_attack}}=5.04$, $p<.001$), suggesting that participants were more sensitive to factual inconsistencies when relying on their own knowledge. No significant difference was observed in Physical and Engineering Sciences ($p=0.42$). These findings suggest that the effectiveness of adversarial explanations is domain-dependent, with fact-heavy fields being more susceptible to persuasive but misleading content.
Table 8: Comparison of trust scores $T$ across task domains between attack and non-attack conditions. For each domain, we report sample size ($n$), mean, variance, and Welch’s t-test results.
Domain Condition $n$ $\bar{T}$ $\mathrm{Var}$ t p
Biological, Medical, and Health Sciences Attack 996 5.10 3.17 3.39 $7.41\times 10^{-4}$
Non-Attack 397 4.76 2.78
Mathematical Problem and Logical Reasoning Attack 826 3.95 3.77 -6.70 $1.41\times 10^{-10}$
Non-Attack 162 4.98 3.04
Social Sciences and Law Attack 994 3.92 3.57 -8.44 $1.38\times 10^{-15}$
Non-Attack 194 5.04 2.75
Physical and Engineering Sciences Attack 986 4.68 3.14 -0.81 $4.17\times 10^{-1}$
Non-Attack 394 4.76 2.85
Politics and History Attack 990 3.82 3.59 -10.31 $1.84\times 10^{-23}$
Non-Attack 394 4.92 3.09
Business and Management Attack 990 5.07 2.77 5.40 $8.75\times 10^{-8}$
Non-Attack 395 4.56 2.36
C.4 Familiarity with AI
Users with little or no familiarity with AI exhibit higher trust in adversarial explanations compared to those with expert-level familiarity ($\bar{T}$: $4.72$ vs. $4.21$, Figure 17). The distributions for novices are right-skewed, with a clear concentration around high trust scores (5–6) and an absence of any score of 1, indicating that less-experienced users are especially vulnerable and tend to accept misleading content at face value. In contrast, AI experts report consistently lower trust in both attack and non-attack scenarios (attack: 4.21, non-attack: 4.59). As familiarity increases, participants assign more low trust scores (1–2) and are also less likely to attribute trust to the AI system itself (Figure 18). This suggests that AI expertise fosters a more critical and cautious evaluation of AI outputs, whereas a lack of familiarity amplifies susceptibility to persuasive but incorrect explanations.
Figure 17: Less-experienced users retain high trust even when explanations are adversarial, while expert users exhibit lower trust and stronger discernment. Distribution of trust scores $T$ across AI familiarity levels under attacks and non-attacks.
User familiarity with AI moderates trust in adversarial explanations, though this effect is uneven across familiarity levels. At the two extremes of the spectrum (i.e., Not familiar at all and Expert-level knowledge) sample sizes are relatively small ($n=29/10$ and $n=145/46$, respectively), and no statistically significant trust difference is observed between attack and non-attack in Table 9.
Users with intermediate levels of AI familiarity constitute the majority of the sample and exhibit a consistent and statistically robust pattern. Participants who are Slightly familiar, Moderately familiar, and Very familiar all report significantly lower trust under attack than under non-attack conditions. Across these groups, trust distributions shift leftward under attack, indicating stronger discernment and reduced susceptibility to misleading explanations.
Table 9: Comparison of trust scores $T$ across AI familiarity levels between attack and non-attack conditions. For each level, we report sample size ($n$), mean, variance, and Welch’s t-test results.
Familiarity Level Condition $n$ $\bar{T}$ $\mathrm{Var}$ t p
Not familiar at all Attack 29 4.72 1.42 0.27 0.790
Non-Attack 10 4.60 1.60
Slightly familiar Attack 466 4.31 3.24 -4.37 $1.75\times 10^{-5}$
Non-Attack 158 4.97 2.60
Moderately familiar Attack 2931 4.44 3.63 -5.38 $8.35\times 10^{-8}$
Non-Attack 981 4.78 2.76
Very familiar Attack 2211 4.46 3.65 -4.44 $9.73\times 10^{-6}$
Non-Attack 741 4.79 2.90
Expert-level knowledge Attack 145 4.21 4.49 -1.16 0.250
Non-Attack 46 4.59 3.36
Figure 18 decomposes user trust by cognitive source across familiarity levels under attack and non-attack conditions. Across all familiarity levels, explanation-based trust remains the dominant source in both conditions. The Not familiar at all group attributes nearly all trust to explanations with a limited sample size.
As AI familiarity increases, reliance on explanations decreases modestly, while attributions to prior knowledge and trust in AI become more pronounced. This shift is especially evident among users with moderate to high familiarity, who also exhibit lower mean trust under attack. These users appear more likely to cross-check explanations against their own understanding, resulting in reduced acceptance of adversarial content. The trust-in-AI attributions remain relatively stable across familiarity levels and conditions, suggesting that system-level trust reflects a persistent user attitude rather than sensitivity to explanation correctness.
Figure 18: Mean trust by cognitive sources (top) and proportions (bottom) across users’ familiarity with AI under attacks and non-attacks.
C.5 Education Level
Education level systematically moderates users’ reliance to adversarial explanations. As shown in Table 10 and Figure 7, trust scores are consistently lower under attack across all education levels, the magnitude of this reduction varies substantially.
Participants with lower educational levels exhibit relatively small trust differences between the attack and non-attack conditions (high school: $\Delta T=0.26$; college: $\Delta T=0.19$), indicating limited sensitivity to the correctness of the explanation. In contrast, users holding bachelor’s and master’s degrees demonstrate markedly larger trust gaps ($\Delta T=0.53$ and $0.57$, respectively), with highly significant differences ($p\ll.001$), suggesting stronger resistance to adversarial explanations and greater reliance on correctness cues. Although users with doctoral or professional degrees also show a large numerical trust difference ($\Delta T=0.56$), this effect is statistically weaker ($p=0.022$), likely due to a smaller sample size and higher within-group variance. Overall, these results indicate that higher educational attainment is associated with increased resistance to adversarial explanations.
Table 10: Comparison of trust scores $T$ across education levels between attack and non-attack conditions. We report sample size ($n$), mean trust, variance, trust difference ($\Delta T=\bar{T}_{\text{non\_attack}}-\bar{T}_{\text{attack}}$), and Welch’s two-sample t-test results.
Education Level Group $n$ $\bar{T}$ Var $\Delta T$ $t$ $p$
High school diploma or GED Attack 1090 4.37 3.13 0.26 -2.65 $8.22\times 10^{-3}$
Non-Attack 370 4.63 2.50
College or associate degree Attack 2034 4.53 3.51 0.19 -2.43 $1.54\times 10^{-2}$
Non-Attack 690 4.72 2.88
Bachelor’s degree Attack 1445 4.40 3.90 0.53 -5.77 $1.07\times 10^{-8}$
Non-Attack 482 4.93 2.75
Master’s degree Attack 1009 4.38 3.89 0.57 -4.92 $1.10\times 10^{-6}$
Non-Attack 325 4.95 3.05
Doctoral or professional degree Attack 204 4.31 3.89 0.56 -2.31 $2.23\times 10^{-2}$
Non-Attack 69 4.87 2.67
Figure 19 further reveals how education level reshapes the cognitive basis of trust under both attack and non-attack conditions. Across all education levels, explanation-based trust remains the dominant source, but its relative importance decreases with higher educational attainment. Users with a high school diploma or some college rely primarily on explanation plausibility, with limited attribution to prior knowledge. In contrast, participants holding bachelor’s and master’s degrees exhibit a pronounced shift toward prior-knowledge-based trust, indicating greater reliance on internal verification when assessing explanation correctness.
This redistribution of cognitive sources aligns with the observed trust reductions under attack among higher-education users. As reliance on prior knowledge increases, adversarial explanations are more likely to conflict with users’ internal expectations, resulting in lower trust. Conversely, users with less formal education exhibit more trust attribution patterns dominated by explanation-based reasoning, making them more susceptible to persuasive but incorrect explanations.
Figure 19: Mean trust by cognitive sources (top) and proportions (bottom) across education levels under attacks and non-attacks.
C.6 Age
Age systematically moderates users’ resistance to adversarial explanations. As shown in Table 11, the youngest users (18–24) exhibit only a small and statistically non-significant trust difference under attack ($\Delta T=0.13$, $p=0.056$), indicating limited sensitivity to misleading explanations. In contrast, all older age groups show significant trust reductions under attack, with substantially larger trust gaps. Older users aged 65 and above show the largest numerical trust gap ($\Delta T=1.02$, $p<0.001$), which suggests older users tend to demonstrate greater skepticism toward AI explanations under adversarial conditions. Users aged 25–34 and 35–44 display moderate resistance ($\Delta T=0.49$ and $0.45$, respectively; $p<10^{-3}$), the middle-aged users (45–54 and 55–64) exhibit the stronger effects ($\Delta T=0.58$ in both groups; $p<10^{-4}$), indicating pronounced discernment when explanations are incorrect.
Overall, these results suggest that susceptibility to adversarial explanations decreases with age. Younger users maintain higher baseline trust and show limited differentiation between correct and incorrect explanations, whereas older users are more likely to penalize misleading explanations, reflecting stronger evaluative judgment and experience-driven skepticism.
Table 11: Comparison of trust scores $T$ across age groups between attack and non-attack conditions. We report sample size ($n$), mean trust, variance, trust difference, and Welch’s two-sample t-test results.
Age Group Group $n$ $\bar{T}$ Var $\Delta T$ $t$ $p$
18–24 Attack 2418 4.49 3.30 0.13 -1.91 $5.60\times 10^{-2}$
Non-Attack 819 4.62 2.77
25–34 Attack 1125 4.42 3.86 0.49 -4.49 $8.57\times 10^{-6}$
Non-Attack 362 4.91 3.14
35–44 Attack 815 4.11 3.90 0.45 -3.64 $3.03\times 10^{-4}$
Non-Attack 279 4.56 2.78
45–54 Attack 680 4.50 3.66 0.58 -4.75 $2.70\times 10^{-6}$
Non-Attack 231 5.08 2.18
55–64 Attack 599 4.74 3.58 0.58 -4.13 $4.50\times 10^{-5}$
Non-Attack 195 5.32 2.62
65 or older Attack 145 3.92 4.19 1.02 -3.62 $4.50\times 10^{-4}$
Non-Attack 50 4.94 2.55
Figure 20 illustrates how age reshapes both the magnitude and the cognitive basis of trust under attack and non-attack conditions. Across all age groups, explanation-based trust remains the primary source; however, its dominance diminishes with age. Younger users (18–24) rely most heavily on explanations and exhibit similar mean trust under attack and non-attack conditions, indicating limited sensitivity to correctness. Consistent with this pattern, the proportion of prior-knowledge-based trust is lowest among the youngest group. As age increases, users increasingly attribute trust to prior knowledge, accompanied by larger trust reductions under attack. Middle-aged users (35–54) show a pronounced shift away from explanation-based trust toward prior knowledge, aligning with the substantial trust gaps observed in this range. Older users (55–64 and 65+) display the strongest reliance on prior knowledge and the lowest mean trust under attack, suggesting heightened scrutiny when explanations conflict with accumulated experience. Trust attributed directly to the AI system remains relatively stable across age groups and conditions, indicating that system-level trust is less sensitive to age than explanation evaluation.
Figure 20: Mean trust by cognitive sources (top) and proportions (bottom) across ages under attacks and non-attacks.
C.7 Initial Trust Levels
Users’ initial attitudes toward AI strongly condition their response to adversarial explanations. As shown in Table 12, users who initially report high trust in AI exhibit the largest trust reductions under attack. In particular, the “Strongly trust” group shows the greatest trust gap ($\Delta T=1.18$, $p<10^{-3}$), indicating that misleading explanations substantially erode confidence when strong prior trust is violated. The “Somewhat trust” and “Neutral” groups also demonstrate significant, though more moderate, trust drops ($\Delta T=0.31$ and $0.30$, respectively; $p\ll 0.001$).
Users who initially express skepticism toward AI show a different pattern. While the “Somewhat distrust” group still exhibits a significant trust reduction under attack ($\Delta T=0.56$, $p<10^{-6}$), the “Strongly distrust” group shows no significant difference between attack and non-attack conditions ($\Delta T=-0.06$, $p=0.85$). This suggests a floor effect: users who already distrust AI maintain consistently low trust regardless of explanation correctness.
Overall, these results indicate that adversarial explanations are most disruptive for users with moderate to high initial trust, while users with entrenched distrust remain largely unaffected. This highlights initial trust as a key moderating factor in adversarial explanation effectiveness.
Table 12: Comparison of trust scores $T$ across users’ initial trust levels between attack and non-attack conditions. We report sample size ($n$), mean trust, variance, trust difference ($\Delta T$), and Welch’s two-sample t-test results.
Initial Trust Level Group $n$ $\bar{T}$ Var $\Delta T$ $t$ $p$
Strongly trust Attack 146 4.60 4.52 1.18 -4.01 $1.28\times 10^{-4}$
Non-Attack 41 5.78 2.28
Somewhat trust Attack 2429 4.80 3.36 0.31 -4.64 $3.85\times 10^{-6}$
Non-Attack 815 5.11 2.57
Neutral Attack 1997 4.34 3.47 0.30 -3.98 $7.33\times 10^{-5}$
Non-Attack 671 4.64 2.78
Somewhat distrust Attack 1078 3.86 3.65 0.56 -5.33 $1.34\times 10^{-7}$
Non-Attack 365 4.42 2.68
Strongly distrust Attack 132 3.63 3.58 -0.06 0.19 0.853
Non-Attack 44 3.57 3.51
Figure 21: Mean trust by cognitive sources (top) and proportions (bottom) across users’ initial trust levels in AI under attacks and non-attacks.
C.8 Long-term Effect
We observe a gradual erosion of trust with repeated task exposure. As shown in Figure 10, average trust decreases as users progress through the task sequence. A linear mixed-effects model (LMM) confirms a significant negative association between task order and trust ($\beta=-0.005$, $p=0.007$, Table 13), indicating that repeated exposure to adversarial explanations cumulatively reduces user trust over time.
To further quantify the cumulative effects of repeated exposure, we analyze the relationship between streak length and the trust score $T$ in the subsequent task using Spearman’s rank correlation test. As shown in Table 14, the length of detected attack streaks is significantly negatively correlated with subsequent trust ($\rho=-0.276$, $p<.001$), indicating that repeated exposure to adversarial explanations progressively erodes user trust. In contrast, benign streak length exhibits a strong positive correlation with subsequent trust ($\rho=0.320$, $p<.001$), suggesting that sustained exposure to correct explanations can restore and reinforce trust over time.
Together with the linear mixed-effects model results on task order, these findings demonstrate that while trust resets at the task level in the short term, it evolves cumulatively over longer horizons through repeated adversarial or benign exposure.
Table 13: LMM results for long-term trust dynamics. The model predicts trust scores $T$ as a function of task order, with random intercepts for participants to account for repeated measures. Significance: $^{*}p<.1$, $^{**}p<.01$, $^{***}p<.001$.
Effect $\boldsymbol{\beta}$ STD $p$
Intercept $4.58^{***}$ 0.08 <.001<.001
Task Order $005^{**} 0.0$02 0.007
Group Var = 0.888
Table 14: Spearman rank correlation analysis between streak length and trust in the subsequent task.
Streak Type $\boldsymbol{\rho}$ $p$-value
Attack streak 0.276-0.276 $10^{-100}$
Non-attack streak $0.320$ $10^{-140}$